# SoliReward: Mitigating Susceptibility to Reward Hacking and Annotation Noise in Video Generation Reward Models

### v1

@@ Abstract
Post-training alignment of video generation models with human preferences is a critical goal. Developing effective Reward Models (RMs) for this process faces significant methodological hurdles. Current data collection paradigms, reliant on in-prompt pairwise annotations, suffer from labeling noise. Concurrently, the architectural design of VLM-based RMs, particularly their output mechanisms, remains underexplored. Furthermore, RM is susceptible to reward hacking in post-training. To mitigate these limitations, we propose SoliReward, a systematic framework for video RM training. Our framework first sources high-quality, cost-efficient data via single-item binary annotations, then constructs preference pairs using a cross-prompt pairing strategy. Architecturally, we employ a Hierarchical Progressive Query Attention mechanism to enhance feature aggregation. Finally, we introduce a modified BT loss that explicitly accommodates win-tie scenarios. This approach regularizes the RM’s score distribution for positive samples, providing more nuanced preference signals to alleviate over-focus on a small number of top-scoring samples. Our approach is validated on benchmarks evaluating physical plausibility, subject deformity, and semantic alignment, demonstrating improvements in direct RM evaluation metrics and in the efficacy of post-training on video generation models. Code and benchmark will be publicly available.
@@ 1 Introduction
Recent advancements in video generation models, such as Sora 2 [openai_sora2], Veo 3 [google2025veo], and Seedance [gao2025seedance], have exhibited remarkable capabilities in synthesizing high-fidelity, temporally coherent visual content. Behind these achievements, post-training alignment techniques, analogous to Reinforcement Learning from Human Feedback (RLHF) [ouyang2022training] in Large Language Models (LLMs), are becoming a key driving force for enhancing model performance and correcting physical implausibility, visual artifacts and instruction following. Currently, methods such as DanceGRPO [xue2025dancegrpo] and LongChat-Video [team2025longcat] have begun to utilize flow-based GRPO [liu2025flow] methodologies to align video models.
However, the effectiveness of these alignment techniques relies heavily on the capabilities of their core component: the Reward Model (RM). RM is trained to quantify human preferences. However, constructing an RM that can accurately capture complex video qualities faces critical issues. These include data annotation noise and inconsistency, the vulnerability to reward hacking, and the underexplored design space for VLM-based RM architectures. On the data front, the dominant paradigm, pairwise preference learning [liu2025improving], trains on video pairs $(y_{w},y_{l})$ where $y_{w}$ is preferred over $y_{l}$, typically using the Bradley-Terry (BT) loss [bradley1952rank]. While effective, this “in-prompt” annotation (pairing videos from the same prompt) struggles with pairs of comparable quality, where annotator ambiguity introduces substantial label noise. An alternative, point-wise scoring (e.g., using a five-level scale) [liang2024rich], avoids direct comparison but introduces high inter-annotator disagreement due to the subjective and ambiguous nature of intermediate scores. This noise in the preference data fundamentally degrades the RM’s ability to accurately rank video quality. Beyond data noise, RM is highly susceptible to reward hacking in post-training, where the learned proxy objective deviates from intended human preferences [skalse2022defining]. Finally, the architecture used to extract this scalar reward is often insufficiently expressive, leading to a collapse of the reward where scores cluster together. Common methods, such as using the last token’s embedding [zhao2025swift], relying on the output of a dedicated special token [liu2025improving], or using the probability of a “yes/no” token [wu2025rewarddance], may fail to capture the full spectrum of information, from low-level artifacts to high-level semantic alignment, encoded across the model’s layers.
@@
Furthermore, we propose the BT with Win-Tie (BT-WT) loss to alleviate the critical issue of reward hacking [skalse2022defining]. By augmenting standard win-lose pairs with ‘win-tie’ pairs (e.g., pairing two ‘Pass’ samples), it explicitly penalizes score variance within the positive set, forcing the RM to map all high-quality samples to a compact manifold and thus mitigating reward spikes that lead to hacking.
Finally, to address the architectural limitations identified in extracting a reward signal from a VLM [bai2025qwen2] backbone, we move beyond simplistic pooling strategies. We propose the Hierarchical Progressive Query Attention (HPQA), a novel architecture that explicitly aggregates and refines features from multiple transformer layers, fusing low-level visual fidelity with high-level semantic understanding for a more robust reward signal. Our contributions are as follows:
@@ 2.1 Data Construction for Reward Model Training
Pairwise Preference Annotation. The dominant paradigm in RLHF [ouyang2022training] and recent video RMs [kirstain2023pick, ma2025hpsv3] involves querying preferences between two generations from the same prompt. This naturally aligns with BT objectives [bradley1952rank], which excel at modeling relative utility. However, in-prompt annotation suffers from severe decisional ambiguity when candidates exhibit similar quality levels. It forces subjective tie-breaking, injecting noticeable noise. While some efforts introduce “tie” options [liu2025improving], defining standardized thresholds for neutrality remains difficult and does not fully resolve the underlying ambiguity.
Point-wise Scoring Annotation. Alternatively, discrete multi-level Likert scales (e.g., VideoScore [he2024videoscore]) could introduce high-variance labeling noise, particularly at subjective decision boundaries. As human preference is intrinsically continuous [you2025teaching], forcing rigid discrete labels onto borderline samples (e.g., ‘3’ vs. ‘4’ for similar inputs) discards uncertainty and creates artificial quality gaps. This limitation is underscored by VideoScore [he2024videoscore], which reports critically low Inter-Annotator Agreement (IAA) (e.g., Fleiss’ $\kappa<0.1$ [fleiss1973equivalence] in Trial 2). We attribute this poor agreement to these boundary inconsistencies.
Binary Annotation. Recent work such as VisionReward [xu2024visionreward] demonstrates that simplifying tasks to binary checklists achieves annotation consistency ($\sim 89\%$ agreement). However, its BT-objective is confined to learning linear weights, degenerating the reward into discrete scores and lacking the ranking granularity of more general RMs [liu2025improving]. Our work bridges this critical gap.
@@ 2.2 Architecture of Reward Model
Vision Language Models (VLMs) [bai2025qwen2] have emerged as the primary architectural choice for reward models, such as  [ma2025hpsv3, liu2025improving, wu2025rewarddance, xu2024visionreward, he2024videoscore]. Commonly, a Vision Transformer processes visual inputs, mapping them into the language embedding space. A language backbone then uniformly processes these visual representations alongside textual inputs. However, the methodology for extracting a scalar reward score from the VLM’s output representations remains an open question. Three primary approaches are commonly employed to derive this score:
Linear Head [zhao2025swift]. This method utilizes the embedding of the last token from the VLM’s final hidden state, which is then fed into a Multi-Layer Perceptron (MLP) to produce the scalar score.
Special Token Head [liu2025improving]. A dedicated special token (e.g., <VQ>, <MQ>, <TA> in VideoAlign) is prepended to the input sequence. The score is constructed by selectively isolating and extracting the logits from these pre-defined special token positions, with each token’s output mapping directly to a specific reward dimension.
Token Probability [wu2025rewarddance]. This method reframes the task as a prompted binary decision. The VLM is instructed to evaluate whether the visual input meets specific criteria and to respond with a binary answer (e.g., “yes” or “no”). The reward is then extracted from the probability/logits assigned to the “yes” token in the model’s output.
@@ 2.3 Post-Training on Visual Generation
Group Relative Policy Optimization. Adapted for visual generation, GRPO-based online algorithms [xue2025dancegrpo, liu2025flow, li2025branchgrpo, luo2025sample] addresses the instability of standard policy gradient methods. They utilize online feedback from reward model as guidance for optimization. By introducing Brownian motion, the ordinary differential equation (ODE) inherent in flow matching [lipman2022flow] is transformed into a stochastic differential equation (SDE). This infusion of stochasticity enhances the model’s exploration within the generation space.
Direct Preference Optimization. DPO [liu2025videodpo, liu2025improving, wu2025densedpo, wallace2024diffusion] offers an alternative to computationally intensive policy gradient methods by recasting the constrained RL objective as a classification problem. This reformulation simplifies the process, yielding a more stable training objective. RM can be leveraged to construct preference pairs for DPO training.
Reward Feedback Learning. ReFL [xu2023imagereward, clark2023directly, shen2025directly] is an alignment technique that utilizes a differentiable reward model, employing the reward signal directly as the loss function. This mechanism allows for end-to-end optimization by backpropagating gradients from the reward model directly through the generative model’s parameters, thus avoiding the complexities of GRPO.
@@ 3.1 Data Annotation
The efficacy of an RM depends heavily on preference data quality, yet standard pairwise and point-wise methods suffer from high annotator ambiguity and disagreement [he2024videoscore, liu2025improving]. To mitigate this, we design a single-item binary annotation. By evaluating single videos against objective criteria (Pass/Fail), we diminish label noise and improve IAA, as shown in Sec. 4.1.
@@ 3.2 Cross-Prompt Pairing Strategy
Crucially, this formulation allows us to implement a cross-prompt pairing strategy. The Bradley-Terry model, upon which our loss is based, does not theoretically require pairs to originate from the same prompt [sun2024rethinking]. By pairing “Pass” and “Fail” samples from different prompts, we create a large-scale, diverse dataset that forces the RM to learn generalized representations of quality rather than simple in-prompt relative rankings.
@@ 3.3 Bradley-Terry with Win-Tie
However, during data annotation, annotators may deem two samples to have near-identical quality (i.e., ties, $y_{i}\sim y_{j}$). The vanilla BT loss discards these tie samples, reducing data efficiency. Furthermore, this win-lose only training scheme, particularly when annotating on specific dimensions, can induce reward hacking during subsequent model post-training [skalse2022defining]. Training exclusively on win-lose pairs seeks to maximize the reward margin $r_{\theta}(y_{i})-r_{\theta}(y_{j})$ in Eq. 1. Crucially, this objective imposes no constraints on the reward distribution within the set of positive samples.
@@
In contrast, VideoAlign [liu2025improving] retains all tie samples and employs a loss function that operates on A-wins, B-wins, or a tie (including both win-tie and lose-tie pairs). We argue this strategy is unsuitable for our annotation scenario. Given two independently labeled negative samples (e.g., for deformity), we cannot assert their degree of deformity is equivalent. Erroneously pairing them as a tie would diminish the discriminative capacity of RM and degrade model performance.
@@ 3.4 Hierarchical Progressive Query Attention
A primary advantage of this mechanism is its ability to create a rich feature representation by explicitly aggregating information hierarchically. This approach is conceptually grounded in findings from linguistic analysis in LLMs [vig2019analyzing], which reveals that transformer layers exhibit functional specialization. For instance, their analysis indicates that attention aligns most strongly with syntactic dependency relations in the middle layers , while the deepest layers are most effective at capturing distant relationships. Our mechanism is designed to explicitly harness this hierarchical specialization. The progressive refinement Eq. 4 allows the model to synthesize a query that explicitly bridges these different semantic levels, fusing low-level fidelity with high-level abstraction. Furthermore, the residual connection ensures that this hierarchical information augments rather than replaces the representation from the final layer.
Table 1: Inter-Annotator Agreement (IAA) analysis across 3 annotators. We report Krippendorff’s $\alpha$ [ford2004content], Fleiss’s $\kappa$ [fleiss1973equivalence] and raw agreement for our binary single-item task and pair-wise comparison task. Single-item is more consistent than pairwise annotation.
@@ 4 Experiments
In this section, we empirically evaluate our reward model on both in-domain and out-of-domain accuracy. Additionally, evaluations are demonstrated in post-training on text-to-video (T2V) task. We use the InternVL3 [zhu2025internvl3] series as the reward model backbone and HunyuanVideo [kong2024hunyuanvideo] for post-training verification.
@@ 4.1 Data Annotation and Pairing Strategy
To validate the reliability of our single-item binary annotation approach against traditional pairwise comparisons, we conduct a formal Inter-Annotator Agreement (IAA) analysis, tasking 5 annotators to label a subset of data using both methods. In Tab. 1, we report Krippendorff’s $\alpha$ [ford2004content], Fleiss’s $\kappa$ [fleiss1973equivalence], and the raw percentage agreement. Our single-item task achieves a Moderate agreement, registering $\alpha=0.4939$, $\kappa=0.4925$, and a strong raw agreement of 77.33%. These results are substantially higher than those from the pairwise comparison task, which only reached Fair agreement ($\alpha=0.3516$, $\kappa=0.3494$) with a 54.67% raw agreement. It confirms that our point-wise annotation framework yields more consistent and reliable labels from annotators compared to the conventional pairwise approach.
@@ 4.2.1 Evaluation of Accuracy
Task Approach RM ACC ID OOD Phy & Deform VideoScore [he2024videoscore] 62.30 67.09 VideoScore-v1.1 [he2024videoscore] 60.20 68.65 LiFT [wang2024lift] 52.60 61.11 VisionReward [xu2024visionreward] 58.45 50.75 VideoPhy [bansal2024videophy] 67.35 65.10 UnifiedReward [wang2025unified] 49.90 44.18 VideoAlign [liu2025improving] 54.40 71.60 Ours 78.48 80.08 TA VideoScore [he2024videoscore] 52.75 54.05 VideoScore-v1.1 [he2024videoscore] 51.20 53.36 LiFT [wang2024lift] 1.75 7.40 VisionReward [xu2024visionreward] 30.75 22.24 VideoPhy [bansal2024videophy] 54.85 60.52 UnifiedReward [wang2025unified] 16.45 10.63 VideoAlign [liu2025improving] 49.50 49.14 Ours 79.02 60.25
@@
We benchmark the accuracy of SoliReward against representative baselines on both in-domain (ID) and out-of-domain (OOD) data. ID accuracy is measured on a held-out partition of our training dataset, which is unseen during training process. For OOD evaluation, we employ videos generated by state-of-the-art (SOTA) models (including Wan2.1 [wan2025wan], Wan2.2 [wan2025wan], Veo 3 [google2025veo] and Seedance 1.0 [gao2025seedance]) from a curated prompt set. Video quality is then assessed by human annotators across three criteria: Physical Plausibility, Subject Deformity (Phy & Deform), and Semantic Alignment (TA). Evaluation on OOD dataset is critical [wu2025rewarddance] because ID accuracy is an insufficient indicator of downstream task performance. OOD accuracy, which assesses generalization capability on shifted distributions, serves as a more robust predictor of a model’s final efficacy. Consequently, OOD accuracy emerges as a more critical metric for evaluating a model’s true utility.

### v2

@@ Abstract
Post-training alignment of video generation models with human preferences is a critical goal. Developing effective Reward Models (RMs) for this process faces significant methodological hurdles. Current data collection paradigms, reliant on in-prompt pairwise annotations, suffer from labeling noise. Concurrently, the architectural design of VLM-based RMs, particularly their output mechanisms, remains underexplored. Furthermore, RM is susceptible to reward hacking in post-training. To mitigate these limitations, we propose SoliReward, a systematic framework for video RM training. Our framework first sources high-quality, cost-efficient data via single-item binary annotations, then constructs preference pairs using a cross-prompt pairing strategy. Architecturally, we employ a Hierarchical Progressive Query Attention mechanism to enhance feature aggregation. Finally, we introduce a modified BT loss that explicitly accommodates win-tie scenarios. This approach regularizes the RM’s score distribution for positive samples, providing more nuanced preference signals to alleviate over-focus on a small number of top-scoring samples. Our approach is validated on benchmarks evaluating physical plausibility, subject deformity, and semantic alignment, demonstrating improvements in direct RM evaluation metrics and in the efficacy of post-training on video generation models. Code and benchmark are available at https://github.com/lian700/SoliReward
@@ 1 Introduction
Recent advancements in video generation models, such as Sora 2 [21], Veo 3 [8], and Seedance [7], have exhibited remarkable capabilities in synthesizing high-fidelity, temporally coherent visual content. Behind these achievements, post-training alignment techniques, analogous to Reinforcement Learning from Human Feedback (RLHF) [22] in Large Language Models (LLMs), are becoming a key driving force for enhancing model performance and correcting physical implausibility, visual artifacts and instruction following. Currently, methods such as DanceGRPO [36] and LongChat-Video [26] have begun to utilize flow-based GRPO [16] methodologies to align video models.
However, the effectiveness of these alignment techniques relies heavily on the capabilities of their core component: the Reward Model (RM). RM is trained to quantify human preferences. However, constructing an RM that can accurately capture complex video qualities faces critical issues. These include data annotation noise and inconsistency, the vulnerability to reward hacking, and the underexplored design space for VLM-based RM architectures. On the data front, the dominant paradigm, pairwise preference learning [17], trains on video pairs $(y_{w},y_{l})$ where $y_{w}$ is preferred over $y_{l}$, typically using the Bradley-Terry (BT) loss [3]. While effective, this “in-prompt” annotation (pairing videos from the same prompt) struggles with pairs of comparable quality, where annotator ambiguity introduces substantial label noise. An alternative, point-wise scoring (e.g., using a five-level scale) [14], avoids direct comparison but introduces high inter-annotator disagreement due to the subjective and ambiguous nature of intermediate scores. This noise in the preference data fundamentally degrades the RM’s ability to accurately rank video quality. Beyond data noise, RM is highly susceptible to reward hacking in post-training, where the learned proxy objective deviates from intended human preferences [24]. Finally, the architecture used to extract this scalar reward is often insufficiently expressive, leading to a collapse of the reward where scores cluster together. Common methods, such as using the last token’s embedding [38], relying on the output of a dedicated special token [17], or using the probability of a “yes/no” token [32], may fail to capture the full spectrum of information, from low-level artifacts to high-level semantic alignment, encoded across the model’s layers.
@@
Furthermore, we propose the BT with Win-Tie (BT-WT) loss to alleviate the critical issue of reward hacking [24]. By augmenting standard win-lose pairs with ‘win-tie’ pairs (e.g., pairing two ‘Pass’ samples), it explicitly penalizes score variance within the positive set, forcing the RM to map all high-quality samples to a compact manifold and thus mitigating reward spikes that lead to hacking.
Finally, to address the architectural limitations identified in extracting a reward signal from a VLM [1] backbone, we move beyond simplistic pooling strategies. We propose the Hierarchical Progressive Query Attention (HPQA), a novel architecture that explicitly aggregates and refines features from multiple transformer layers, fusing low-level visual fidelity with high-level semantic understanding for a more robust reward signal. Our contributions are as follows:
@@ 2.1 Data Construction for Reward Model Training
Pairwise Preference Annotation. The dominant paradigm in RLHF [22] and recent video RMs [11, 20] involves querying preferences between two generations from the same prompt. This naturally aligns with BT objectives [3], which excel at modeling relative utility. However, in-prompt annotation suffers from severe decisional ambiguity when candidates exhibit similar quality levels. It forces subjective tie-breaking, injecting noticeable noise. While some efforts introduce “tie” options [17], defining standardized thresholds for neutrality remains difficult and does not fully resolve the underlying ambiguity.
Point-wise Scoring Annotation. Alternatively, discrete multi-level Likert scales (e.g., VideoScore [9]) could introduce high-variance labeling noise, particularly at subjective decision boundaries. As human preference is intrinsically continuous [37], forcing rigid discrete labels onto borderline samples (e.g., ‘3’ vs. ‘4’ for similar inputs) discards uncertainty and creates artificial quality gaps. This limitation is underscored by VideoScore [9], which reports critically low Inter-Annotator Agreement (IAA) (e.g., Fleiss’ $\kappa<0.1$ [5] in Trial 2). We attribute this poor agreement to these boundary inconsistencies.
Binary Annotation. Recent work such as VisionReward [34] demonstrates that simplifying tasks to binary checklists achieves annotation consistency ($\sim 89\%$ agreement). However, its BT-objective is confined to learning linear weights, degenerating the reward into discrete scores and lacking the ranking granularity of more general RMs [17]. Our work bridges this critical gap.
@@ 2.2 Architecture of Reward Model
Vision Language Models (VLMs) [1] have emerged as the primary architectural choice for reward models, such as  [20, 17, 32, 34, 9]. Commonly, a Vision Transformer processes visual inputs, mapping them into the language embedding space. A language backbone then uniformly processes these visual representations alongside textual inputs. However, the methodology for extracting a scalar reward score from the VLM’s output representations remains an open question. Three primary approaches are commonly employed to derive this score:
Linear Head [38]. This method utilizes the embedding of the last token from the VLM’s final hidden state, which is then fed into a Multi-Layer Perceptron (MLP) to produce the scalar score.
Special Token Head [17]. A dedicated special token (e.g., <VQ>, <MQ>, <TA> in VideoAlign) is prepended to the input sequence. The score is constructed by selectively isolating and extracting the logits from these pre-defined special token positions, with each token’s output mapping directly to a specific reward dimension.
Token Probability [32]. This method reframes the task as a prompted binary decision. The VLM is instructed to evaluate whether the visual input meets specific criteria and to respond with a binary answer (e.g., “yes” or “no”). The reward is then extracted from the probability/logits assigned to the “yes” token in the model’s output.
@@ 2.3 Post-Training on Visual Generation
Group Relative Policy Optimization. Adapted for visual generation, GRPO-based online algorithms [36, 16, 13, 19] addresses the instability of standard policy gradient methods. They utilize online feedback from reward model as guidance for optimization. By introducing Brownian motion, the ordinary differential equation (ODE) inherent in flow matching [15] is transformed into a stochastic differential equation (SDE). This infusion of stochasticity enhances the model’s exploration within the generation space.
Direct Preference Optimization. DPO [18, 17, 33, 28] offers an alternative to computationally intensive policy gradient methods by recasting the constrained RL objective as a classification problem. This reformulation simplifies the process, yielding a more stable training objective. RM can be leveraged to construct preference pairs for DPO training.
Reward Feedback Learning. ReFL [35, 4, 23] is an alignment technique that utilizes a differentiable reward model, employing the reward signal directly as the loss function. This mechanism allows for end-to-end optimization by backpropagating gradients from the reward model directly through the generative model’s parameters, thus avoiding the complexities of GRPO.
@@ 3.1 Data Annotation
The efficacy of an RM depends heavily on preference data quality, yet standard pairwise and point-wise methods suffer from high annotator ambiguity and disagreement [9, 17]. To mitigate this, we design a single-item binary annotation. By evaluating single videos against objective criteria (Pass/Fail), we diminish label noise and improve IAA, as shown in Sec. 4.1.
@@ 3.2 Cross-Prompt Pairing Strategy
Crucially, this formulation allows us to implement a cross-prompt pairing strategy. The Bradley-Terry model, upon which our loss is based, does not theoretically require pairs to originate from the same prompt [25]. By pairing “Pass” and “Fail” samples from different prompts, we create a large-scale, diverse dataset that forces the RM to learn generalized representations of quality rather than simple in-prompt relative rankings.
@@ 3.3 Bradley-Terry with Win-Tie
However, during data annotation, annotators may deem two samples to have near-identical quality (i.e., ties, $y_{i}\sim y_{j}$). The vanilla BT loss discards these tie samples, reducing data efficiency. Furthermore, this win-lose only training scheme, particularly when annotating on specific dimensions, can induce reward hacking during subsequent model post-training [24]. Training exclusively on win-lose pairs seeks to maximize the reward margin $r_{\theta}(y_{i})-r_{\theta}(y_{j})$ in Eq. 1. Crucially, this objective imposes no constraints on the reward distribution within the set of positive samples.
@@
In contrast, VideoAlign [17] retains all tie samples and employs a loss function that operates on A-wins, B-wins, or a tie (including both win-tie and lose-tie pairs). We argue this strategy is unsuitable for our annotation scenario. Given two independently labeled negative samples (e.g., for deformity), we cannot assert their degree of deformity is equivalent. Erroneously pairing them as a tie would diminish the discriminative capacity of RM and degrade model performance.
@@ 3.4 Hierarchical Progressive Query Attention
A primary advantage of this mechanism is its ability to create a rich feature representation by explicitly aggregating information hierarchically. This approach is conceptually grounded in findings from linguistic analysis in LLMs [27], which reveals that transformer layers exhibit functional specialization. For instance, their analysis indicates that attention aligns most strongly with syntactic dependency relations in the middle layers , while the deepest layers are most effective at capturing distant relationships. Our mechanism is designed to explicitly harness this hierarchical specialization. The progressive refinement Eq. 4 allows the model to synthesize a query that explicitly bridges these different semantic levels, fusing low-level fidelity with high-level abstraction. Furthermore, the residual connection ensures that this hierarchical information augments rather than replaces the representation from the final layer.
Table 1: Inter-Annotator Agreement (IAA) analysis across 5 annotators. We report Krippendorff’s $\alpha$ [6], Fleiss’s $\kappa$ [5] and raw agreement for our binary single-item task and pair-wise comparison task. Single-item is more consistent than pairwise annotation.
@@ 4 Experiments
In this section, we empirically evaluate our reward model on both in-domain and out-of-domain accuracy. Additionally, evaluations are demonstrated in post-training on text-to-video (T2V) task. We use the InternVL3 [40] series as the reward model backbone and HunyuanVideo [12] for post-training verification.
@@ 4.1 Data Annotation and Pairing Strategy
To validate the reliability of our single-item binary annotation approach against traditional pairwise comparisons, we conduct a formal Inter-Annotator Agreement (IAA) analysis, tasking 5 annotators to label a subset of data using both methods. In Tab. 1, we report Krippendorff’s $\alpha$ [6], Fleiss’s $\kappa$ [5], and the raw percentage agreement. Our single-item task achieves a Moderate agreement, registering $\alpha=0.4939$, $\kappa=0.4925$, and a strong raw agreement of 77.33%. These results are substantially higher than those from the pairwise comparison task, which only reached Fair agreement ($\alpha=0.3516$, $\kappa=0.3494$) with a 54.67% raw agreement. It confirms that our point-wise annotation framework yields more consistent and reliable labels from annotators compared to the conventional pairwise approach.
@@ 4.2.1 Evaluation of Accuracy
Task Approach RM ACC ID OOD Phy & Deform VideoScore [9] 62.30 67.09 VideoScore-v1.1 [9] 60.20 68.65 LiFT [30] 52.60 61.11 VisionReward [34] 58.45 50.75 VideoPhy [2] 67.35 65.10 UnifiedReward [31] 49.90 44.18 VideoAlign [17] 54.40 71.60 Ours 78.48 80.08 TA VideoScore [9] 52.75 54.05 VideoScore-v1.1 [9] 51.20 53.36 LiFT [30] 1.75 7.40 VisionReward [34] 30.75 22.24 VideoPhy [2] 54.85 60.52 UnifiedReward [31] 16.45 10.63 VideoAlign [17] 49.50 49.14 Ours 79.02 60.25
@@
We benchmark the accuracy of SoliReward against representative baselines on both in-domain (ID) and out-of-domain (OOD) data. ID accuracy is measured on a held-out partition of our training dataset, which is unseen during training process. For OOD evaluation, we employ videos generated by state-of-the-art (SOTA) models (including Wan2.1 [29], Wan2.2 [29], Veo 3 [8] and Seedance 1.0 [7]) from a curated prompt set. Video quality is then assessed by human annotators across three criteria: Physical Plausibility, Subject Deformity (Phy & Deform), and Semantic Alignment (TA). Evaluation on OOD dataset is critical [32] because ID accuracy is an insufficient indicator of downstream task performance. OOD accuracy, which assesses generalization capability on shifted distributions, serves as a more robust predictor of a model’s final efficacy. Consequently, OOD accuracy emerges as a more critical metric for evaluating a model’s true utility.